OpenStudy (anonymous):

solve 1/x-5/2x^2 =8/5x Please!

OpenStudy (tyteen4a03):

\(\frac{1}{x} - \frac{5}{2x^2} = \frac{8}{5x}\). Hint: Subtract 8/5x from both sides, then find the LCM of x, 2x^2 and 5x

OpenStudy (anonymous):

help me on my question please anyone

OpenStudy (tyteen4a03):

@HorseCrazyGirlForever I gave out an hint. Which part are you stuck on?

OpenStudy (anonymous):

The whole thing.